# Federated Class-Incremental Learning with Hierarchical Generative Prototypes

### v3

@@ 4.2 Results
We assess the performance of all methods using two widely adopted metrics in FCIL literature: Final Average Accuracy (FAA) and Average Incremental Accuracy (AIA). For the mathematical definitions of these measures, we refer the reader to the supplementary material.
@@
Our proposal is evaluated alongside ten competitors, following the methodology outlined in (Guo et al. 2024). Among these, five were originally designed for Class-Incremental Learning (CIL), one for Federated Learning (FL), and the remainder four for Federated Class-Incremental Learning (FCIL). Following common practice, we adapt CIL approaches for FCIL by aggregating local models with FedAvg (McMahan et al. 2017). Instead, the FL methodology (CCVR) is adapted for FCIL by enabling it to leverage all artifacts produced across tasks.
From CIL, we include two regularization-based techniques (EWC (Kirkpatrick et al. 2017), LwF (Li and Hoiem 2017)), one leveraging a rehearsal buffer (iCaRL (Rebuffi et al. 2017)), and two prompting-based approaches (L2P (Wang et al. 2022b), CODA-Prompt (Smith et al. 2023)). From FL and FCIL, we select a rebalancing-based approach (CCVR (Luo et al. 2021)), a Parameter-Efficient Fine-Tuning technique (PLoRA (Guo et al. 2024)), two rehearsal-based methods (GLFC (Dong et al. 2022), LGA (Dong et al. 2023)), and one utilizing generative replay (TARGET (Zhang et al. 2023b)). Results for all methods, except HGP, CODA-Prompt, and CCVR, are sourced from previous research (Guo et al. 2024). Finally, we include the upper bound, established by training the same backbone jointly on the global data distribution (i.e., without applying of federated or incremental splitting), referred to as Joint.
@@
In Tables 1 and 2, we present the results of HGP and the other approaches in terms of FAA []delimited-[][$\uparrow$][ ↑ ] and AIA []delimited-[][$\uparrow$][ ↑ ] on the CIFAR-100 and Tiny-ImageNet datasets, respectively. Generally, all methods tested on Tiny-ImageNet (including Joint) exhibit lower performance compared to CIFAR-100, indicating the greater challenge posed by the former dataset. Consistent with previous findings (van de Ven, Tuytelaars, and Tolias 2022; Buzzega et al. 2020), regularization techniques for CIL (EWC, LwF) perform poorly across all imbalance levels, indicating severe forgetting issues during incremental learning. iCaRL is less effective compared to prompting techniques like L2P [^1] For further insights on L2P results, we refer the reader to the supplementary material. and CODA-Prompt on more balanced datasets. However, thanks to the presence of a memory buffer of old samples, it maintains relatively consistent results as label imbalance increases. Even without a buffer, CCVR also maintains stable performance thanks to its effective rebalancing strategy, providing better results than iCaRL. While L2P generally outperforms CODA-Prompt, its performance remains lower than PEFT-based methods designed specifically for FCIL.
Among FCIL methods, TARGET exhibits the lowest performance. Notably, it exhibits significantly higher AIA compared to FAA, suggesting ineffective mitigation of forgetting through its generative replay. GLFC and LGA [^2] Due to computing constraints, we omit the performance of GLFC and LGA on distribution-based label imbalance settings. perform better than iCaRL but worse than prompting techniques, indicating the improved effectiveness of their class-aware gradient compensation compared to relying solely on a memory buffer. However, both methods struggle with highly imbalanced label distributions.
Partition $\beta=0.5$ $\beta=0.1$ $\beta=0.05$ $\alpha=12$ $\alpha=8$ $\alpha=4$
Joint $\boldsymbol{84.3}$
CCVR $56.4$ $(71.7)$ $58.6$ $(72.2)$ $56.3$ $(69.8)$ $55.4$ $(71.6)$ $55.6$ $(71.7)$ $57.2$ $(70.4)$
L2PFL $64.2$ $(66.9)$ $56.3$ $(52.5)$ $51.9$ $(43.2)$ $61.6$ $(58.0)$ $49.4$ $(39.3)$   $8.2$ $(10.2)$
CODA-PFL $72.2$ $(81.1)$ $54.4$ $(64.2)$ $45.6$ $(57.9)$ $64.3$ $(75.1)$ $49.9$ $(62.2)$ $27.1$ $(42.8)$
TARGET   $7.7$ $(20.5)$ $0.5$ $(8.7)$ $0.9$ $(5.3)$   $3.5$ $(17.6)$   $4.1$ $(16.4)$ $0.5$ $(4.8)$
GLFC -- ()(-)( - ) -- ()(-)( - ) -- ()(-)( - ) $35.0$ $(47.9)$ $20.2$ $(37.5)$ $5.1$ $(8.7)$
LGA -- ()(-)( - ) -- ()(-)( - ) -- ()(-)( - ) $37.3$ $(53.2)$ $61.4$ $(67.5)$ $6.5$ $(9.8)$
PLoRA $75.5$ $(81.8)$ $74.3$ $(81.0)$ $74.1$ $(80.7)$ $75.3$ $(81.7)$ $75.2$ $(81.6)$ $73.6$ $(80.1)$
HGP (ours) $\boldsymbol{78.6}$ $(\boldsymbol{84.9})$ $\boldsymbol{78.6}$ $(\boldsymbol{84.8})$ $\boldsymbol{78.1}$ $(\boldsymbol{84.8})$ $\boldsymbol{78.4}$ $(\boldsymbol{84.8})$ $\boldsymbol{78.4}$ $(\boldsymbol{84.8})$ $\boldsymbol{78.2}$ $(\boldsymbol{84.4})$
Gain $\texttt{+}3.1$ $(\texttt{+}3.1)$ $\texttt{+}4.3$ $(\texttt{+}3.8)$ $\texttt{+}4.0$ $(\texttt{+}4.1)$ $\texttt{+}3.1$ $(\texttt{+}3.1)$ $\texttt{+}3.2$ $(\texttt{+}3.2)$ $\texttt{+}4.6$ $(\texttt{+}4.3)$
Table 2: Tiny-ImageNet results. Results in terms of FAA []delimited-[][$\uparrow$][ ↑ ] and AIA []delimited-[][$\uparrow$][ ↑ ]. AIA is reported between parenthesis. Best results are highlighted in bold. Gain refers to the previous SOTA approach. The subscript FL indicates CL methods adapted for FCIL.
PEFT-based methods specifically designed for FCIL (PLoRA, HGP) outperform other methodologies across all settings. The use of prototypes enables them to handle increasing label imbalance more effectively, resulting in the most stable performance among variations in $\alpha$ and $\beta$ in both datasets. HGP significantly outperforms PLoRA in all scenarios by leveraging generative prototypes on the centralized server, demonstrating superior robustness to both incremental and federated biases.
@@ 4.3 Ablation study
In this section, we explore the impact of prompting on Federated Bias, emphasizing the contrast with the traditional approach of fine-tuning the whole network. All experiments are performed on the CIFAR-100 dataset, distributed across $10$ clients under different distribution-based label imbalance settings.
We argue that prompting is a more advantageous approach for adapting pre-trained models in Federated Learning, as it constrains the bias to the last layer instead of distributing it across the whole network. To experimentally prove this claim, we position ourselves at the end of the local training prior to the first synchronization with the server: namely, when each client is trained exclusively on its local distribution. In this scenario, we assess the Federated Bias within the feature space – which we refer to as feature bias for short – by computing the local prototypes for the $10$ clients and measuring their average pairwise Euclidean distance. Figure 3 (left) shows that prompting results in smaller feature bias compared to traditional fine-tuning, suggesting that the obtained features are less biased towards the local distributions. To validate this assertion, we also compute the same metric on the pre-trained model (ViT-B-16 on ImageNet-21k), which yields the smallest bias by design. This further supports our reasoning, as this last network experiences no fine-tuning whatsoever.
On the right-hand side of Figure 3, we show the performance of prompting vs. traditional fine-tuning. In line with recent works (Zhang et al. 2023a; Panos et al. 2023; McDonnell et al. 2024), prompting (Prompt) shows inferior performance compared to fine-tuning, due to its limited plasticity. However, when leveraging Classifier Rebalancing (CR) on both approaches, we observe a reversal in this trend. Notably, simply addressing Federated Bias in the last layer is sufficient for prompting to outperform classical fine-tuning: this suggests that prompting techniques train more biased classifiers while producing less biased features, effectively restricting FB to the final layer.
Prompting CRold CRcur CIFAR-100 Tiny-ImageNet
$44.8$ $38.7$
$51.8$ $43.5$
$81.7$ $70.7$
$77.8$ $69.8$
$88.9$ $78.7$
Table 3: HGP components. Evaluation of the singular components of HGP in terms of FAA []delimited-[][$\uparrow$][ ↑ ].
@@
Impact of Different Components.
We assess the specific contributions of various components of HGP in terms of FAA for both benchmarks under one distribution-based label imbalance setting with $\beta=0.05$. Results of these ablative experiments are summarized in Table 3.
Starting with the full fine-tuning of ViT-B/16 serves as a lower bound. Incorporating our prompting technique yields a notable improvement. It is worth noting that the primary advantage of prompting lies not only in enhancing accuracy but also in confining bias to the final classification layer, thereby facilitating Classifier Rebalancing.
We denote CRcur and CRold as the classifier rebalancing procedure applied to the current task classes and the old tasks classes, respectively. Specifically, CRcur rebalances the classifier with features generated from prototypes belonging to the set of classes of the current task, while CRold employs sampling on the set of classes of the old tasks. The latter performs better than the former, indicating the greater importance of addressing Incremental Bias.
By integrating all components, HGP achieves state-of-the-art performance as reported in Tables 1 and 2.
@@ 5 Relation with prior works
In its most naive form, Federated Learning suggests combining local models at the conclusion of each round through parameters averaging (McMahan et al. 2017). Other approaches introduce regularization techniques during local models’ training to prevent them from drifting too far from the server’s parameter space. For instance, FedProx (Li et al. 2020) and SCAFFOLD (Karimireddy et al. 2020) enforce such regularization. FedDC (Gao et al. 2022) estimates the local parameter shift and employs it as a correction term before aggregating parameters. GradMA (Luo et al. 2023) addresses optimization challenges (i.e., quadratic programming) by rerouting each client update towards optimizing the local problem while maintaining proximity to the server. On a different trajectory, FedProto (Tan et al. 2022) computes prototypes for each class observed by each client, which are then aggregated by the server into global prototypes, serving as target representations for subsequent rounds. Similarly, in (Luo et al. 2021), the authors fit Gaussian distributions for each class and use them to generate an IID dataset of features, subsequently calibrating the server’s classification head with these features. In our work, we leverage hierarchical Gaussian Mixture Models (GMMs) of prototypes and sample from them to rebalance the global classifier.
@@
Class-Incremental Learning (CIL) stands out as one of the most challenging settings within the Continual Learning domain (van de Ven, Tuytelaars, and Tolias 2022). In this scenario, the training process is divided into distinct tasks, each introducing new classes as the training progresses. To address this setting, early methods rely on regularization techniques, establishing checkpoints for previous tasks and aiming to maintain proximity w.r.t. their parameters (Zenke, Poole, and Ganguli 2017; Kirkpatrick et al. 2017) or predictions (Li and Hoiem 2017). Alternatively, rehearsal-based approaches involve storing samples in a limited memory buffer and subsequently replaying them to optimize either the original objective (Robins 1995; Chaudhry et al. 2018) or a different one based on knowledge distillation (Rebuffi et al. 2017; Buzzega et al. 2020). The emergence of pre-trained self-attentive architectures (Dosovitskiy et al. 2020) in the Computer Vision domain has paved the way for significant advancements, particularly with the advent of Parameter-Efficient Fine-Tuning (PEFT) techniques. Notably, recent approaches (Wang et al. 2022b, a; Smith et al. 2023) have harnessed prompting to achieve state-of-the-art performance in CIL, eliminating the need for a buffer to replay old samples. Instead, they employ a prompt pool comprising incrementally learned prompts, utilized to condition the network during the forward pass. In our method, we embrace the prompting paradigm, which not only delivers robust performance but also facilitates efficient communication rounds in terms of exchanged parameters.
@@
The concept of Federated Class-Incremental Learning (FCIL) was initially introduced in (Yoon et al. 2021). In their work, the authors propose FedWeIT, which partitions client-side parameters into task-generic and task-specific components. To mitigate interference between clients, they implement sparse learnable masks to selectively extract relevant knowledge for each client. GLFC (Dong et al. 2022) takes a different approach by combining local buffers with class-aware gradient compensation loss. This strategy helps counteract catastrophic forgetting through rehearsal, while also adjusting the magnitude of gradient updates based on whether input samples belong to new or old classes. Building upon this framework, an enhanced version is introduced by the same authors in the LGA paper (Dong et al. 2023). TARGET (Zhang et al. 2023b) tackles forgetting by training a centralized generator network to produce synthetic data, maintaining a similar behavior to the generator used in previous tasks. The generative network populates a buffer after each task, allowing clients to utilize it for rehearsal. Recent advancements in FCIL involve fine-tuning pre-trained models using PEFT techniques. Fed-CPrompt (Bagwe et al. 2023) introduces a regularization term that encourages local prompts to diverge from global ones, enabling them to learn task-specific features. Meanwhile, PLoRA (Guo et al. 2024) integrates LoRA (Hu et al. 2021) with prototypes, which are aggregated via a re-weighting mechanism on the server side during each communication round. As a result, clients leverage these aggregated prototypes as target representations to govern local training.

### v4

@@ 4.2 Results
We assess the performance of all methods using the widely adopted Final Average Accuracy (FAA). For the formal definition of such measure, we refer the reader to the suppl. material.
@@
Our proposal is evaluated alongside $13$ competitors. Among these, five were originally designed for Class-Incremental Learning (CIL), two for Federated Learning (FL), two for Model Merging, and the remainder four for Federated Class-Incremental Learning (FCIL). Following common practice, we adapt CIL approaches for FCIL by aggregating local models with FedAvg mcmahan2017communication . From CIL, we include two regularization-based techniques (EWC kirkpatrick2017overcoming , LwF li2017learning ) and three prompting-based approaches (L2P wang2022learning , DualPrompt wang2022dualprompt , and CODA-Prompt smith2023coda ). From FL, we select a rebalancing method (CCVR luo2021no ), and one using a prototype-based classifier (FedProto tan2022fedproto ); we include two model merging techniques (FisherAVG matena2022merging and RegMean jin2022dataless ), and four FCIL methods. These leverage Parameter-Efficient Fine-Tuning (PILoRA guo2024federated ), generative replay (TARGET zhang2023target ), regularization prototypes (PIP ma2024pip ), and closed-form merging (LoRM salami2024closed ). Finally, we include the upper bound, where the same ViT-B/16 backbone is trained jointly on the global data distribution (i.e., without federated or incremental splitting), referred to as "Joint".
Table 1: Cifar-100, Imagenet-R and Imagenet-A. Results in terms of FAA []delimited-[][$\uparrow$][ ↑ ]. Best results are highlighted in bold, second-best underlined.
CIFAR-100 Imagenet-R Imagenet-A
Joint $\boldsymbol{92.75}$ $\boldsymbol{84.02}$ $\boldsymbol{54.64}$
Partition $\beta$ $0.5$ $0.1$ $0.05$ $0.5$ $0.1$ $0.05$ $1.0$ $0.5$ $0.2$
EWC $78.46$ $72.42$ $64.51$ $58.93$ $48.15$ $43.68$ $10.86$ $10.07$ $8.89$
LwF $62.87$ $55.56$ $47.09$ $54.03$ $41.02$ $46.07$ $8.89$ $8.89$ $7.90$
FisherAVG $76.10$ $74.43$ $65.31$ $58.68$ $50.82$ $47.33$ $11.59$ $11.06$ $10.14$
RegMean $59.80$ $45.88$ $39.08$ $61.18$ $57.00$ $55.80$ $8.56$ $6.22$ $4.34$
CCVR $79.95$ $75.14$ $65.30$ $70.00$ $62.60$ $60.38$ $\underline{39.50}$ $36.27$ $\underline{35.94}$
L2P $83.88$ $61.54$ $55.00$ $42.08$ $23.85$ $16.98$ $20.14$ $17.31$ $16.85$
DualPrompt $80.49$ $54.31$ $42.43$ $45.89$ $27.34$ $21.76$ $20.91$ $16.50$ $9.02$
CODA-P $82.25$ $61.82$ $46.74$ $61.18$ $36.73$ $25.82$ $18.30$ $14.48$ $7.31$
FedProto $75.79$ $70.02$ $60.55$ $58.52$ $47.30$ $52.93$ $9.87$ $9.22$ $10.01$
TARGET $74.72$ $72.32$ $62.60$ $54.65$ $45.83$ $41.32$ $10.27$ $11.39$ $10.73$
PIP $85.01$ $\underline{81.87}$ $81.40$ $58.74$ $55.52$ $56.64$ $26.90$ $26.38$ $24.69$
PILoRA $76.48$ $75.81$ $74.80$ $53.67$ $51.62$ $49.37$ $19.62$ $18.70$ $20.01$
LoRM $\underline{86.95}$ $81.75$ $\underline{82.76}$ $\underline{72.48}$ $\underline{63.83}$ $\underline{66.45}$ $37.26$ $\underline{36.34}$ $33.11$
HGP (ours) $\boldsymbol{90.39}$ $\boldsymbol{90.20}$ $\boldsymbol{90.16}$ $\boldsymbol{72.64}$ $\boldsymbol{71.93}$ $\boldsymbol{71.58}$ $\boldsymbol{41.61}$ $\boldsymbol{41.01}$ $\boldsymbol{40.62}$
@@
In Table 1, we present the results for CIFAR-100, Imagenet-R, and Imagenet-A. Consistently on the first two datasets, prompting methodologies struggling as the label heterogeneity in the data distribution increases. This aligns with our findings, suggesting that federated bias is more pronounced when learning prompts (see Section 4.3). In contrast, FisherAVG and LwF are steadier, but still exhibit poor performance overall. EWC and FisherAVG surprisingly perform similar to FL methodologies (CCVR, FedProto), only to be surpassed by FCIL-targeted approaches (PILoRA, PIP, TARGET, LoRM, HGP).
ImageNet-A, specifically designed to evaluate model robustness, poses a greater challenge by incorporating samples from ImageNet that are frequently misclassified. This is also evidenced by the lower performance of all methods. Here, classical CL techniques struggle to learn the task effectively and are surpassed by prompting methodologies. Among FL and FCIL approaches, only PIP and PILoRA achieve decent performance, with CCVR and LoRM outperforming them. Finally, thanks to the combination of classifier rebalancing and prompting, HGP consistently delivers the best performance in all settings, while LoRM ranks as the second-best performer on average.
Table 2: EuroSAT, Cars-196 and CUB-200. Results in terms of FAA []delimited-[][$\uparrow$][ ↑ ]. Best results are highlighted in bold, second-best underlined.
EuroSAT Cars-196 CUB-200
Joint $\boldsymbol{98.42}$ $\boldsymbol{85.62}$ $\boldsymbol{86.04}$
Partition $\beta$ $0.5$ $0.1$ $0.05$ $0.5$ $0.1$ $0.05$ $1.0$ $0.5$ $0.2$
EWC $64.12$ $59.30$ $56.52$ $19.55$ $18.02$ $18.29$ $31.46$ $29.60$ $27.89$
LwF $31.91$ $21.26$ $31.42$ $20.84$ $22.72$ $31.76$ $25.25$ $21.11$ $18.54$
FisherAVG $58.84$ $59.94$ $55.86$ $26.03$ $24.60$ $21.58$ $30.45$ $28.39$ $25.06$
RegMean $48.74$ $51.73$ $45.27$ $21.83$ $20.36$ $15.92$ $35.57$ $32.84$ $32.83$
CCVR $64.44$ $57.93$ $62.69$ $38.99$ $37.81$ $35.31$ $62.67$ $59.48$ $56.33$
L2P $40.63$ $51.78$ $45.46$ $35.49$ $31.00$ $20.01$ $56.23$ $47.31$ $38.16$
DualPrompt $62.97$ $52.78$ $55.34$ $34.07$ $26.47$ $21.30$ $60.93$ $55.59$ $44.61$
CODA-P $73.38$ $69.42$ $66.69$ $28.04$ $20.83$ $14.53$ $42.53$ $37.71$ $29.19$
FedProto $58.79$ $62.85$ $64.17$ $26.08$ $24.55$ $22.75$ $30.22$ $28.27$ $26.01$
TARGET $52.74$ $52.74$ $45.11$ $28.65$ $27.20$ $26.13$ $39.30$ $38.40$ $34.79$
PIP $58.12$ $53.67$ $54.16$ $36.02$ $34.33$ $29.99$ $\underline{69.60}$ $\underline{65.46}$ $\underline{60.79}$
PILoRA $48.35$ $32.89$ $31.22$ $37.57$ $37.92$ $36.95$ $61.11$ $60.68$ $60.39$
LoRM $\underline{84.23}$ $\underline{77.26}$ $\underline{81.36}$ $\boldsymbol{54.41}$ $\boldsymbol{51.87}$ $\underline{48.81}$ $64.60$ $63.67$ $60.06$
HGP (ours) $\boldsymbol{87.97}$ $\boldsymbol{88.35}$ $\boldsymbol{85.90}$ $\underline{52.13}$ $\underline{51.48}$ $\boldsymbol{49.95}$ $\boldsymbol{80.23}$ $\boldsymbol{79.71}$ $\boldsymbol{78.58}$
Table 2 compares the same methodologies on EuroSAT, Cars-196, and CUB-200. Among these datasets, Cars-196 is the most impacted by the FCIL setting, as it exhibits the largest performance gap between joint training and the best-performing method. In this scenario, EWC and LwF fail to deliver satisfactory results, whereas FL and FCIL methodologies achieve competitive performance. While rebalancing played a critical role in the previously analyzed datasets, CCVR performs slightly worse than PILoRA in this case, highlighting the significant impact of low-rank adaptation. Here, HGP and LoRM compete for the top spot by a wide margin over the others, showing comparable performance.
For EuroSAT and CUB-200, the accuracy gap between Joint training and FCIL techniques narrows, suggesting a less challenging task. FedProto, CODA-Prompt, and DualPrompt demonstrate superior results on EuroSAT, whereas PIP, PILoRA and CCVR underperform on this dataset but achieve the best results on CUB-200. This trend reversal is likely due to the larger number of classes in CUB-200, which challenges traditional CL methods and highlights the significance of approaches specifically tailored for Federated Learning. LoRM emerges as the clear second-best performer on EuroSAT and struggles on CUB-200, where its performance aligns with that of PIP and PILoRA. HGP consistently delivers the highest performance across both datasets, with a notably large margin on CUB-200.
Table 3: HGP components. Evaluation of the impact of each component of HGP. Presented in terms of FAA []delimited-[][$\uparrow$][ ↑ ].
Prompt CRold CRcur C-100 IN-R CUB-200
$30.58$ $26.42$ $25.70$
$52.29$ $30.28$ $26.74$
$81.91$ $66.70$ $69.51$
$85.43$ $68.88$ $47.80$
$90.16$ $71.58$ $78.58$
@@ 4.3 Ablation study
Impact of Different Components.
We assess the specific contributions of various components of HGP in terms of FAA for three datasets under the most challenging distribution-based label imbalance setting ($\beta=0.05$). Results of these ablative experiments are summarized in Table 3, where the full fine-tuning of ViT-B/16 serves as the lower bound. Incorporating our prompting technique yields a remarkable improvement on CIFAR-100 and Imagenet-R, with a minor gain on CUB-200. However, it is worth noting that the primary advantage of prompting lies not only in enhancing accuracy but also in confining bias to the final classification layer, thereby increasing the impact of Classifier Rebalancing.
We define CRcur and CRold as classifier rebalancing procedures applied to the current task classes and the old task classes, respectively. Specifically, CRcur rebalances only the classification head for the current task by utilizing generated features from prototypes corresponding to the current classes. Conversely, CRold focuses on rebalancing the classification heads for previous tasks by sampling from the prototypes of the old classes. On average, both approaches perform comparably; however, CRold demonstrates superior performance on CUB-200, highlighting the importance of addressing Incremental Bias in this dataset. By integrating all components, HGP demonstrates the critical contribution of each, achieving state-of-the-art performance as shown in Tables 1 and 2.
@@
In this section, we explore the impact of prompting on Federated Bias, emphasizing the contrast with the traditional approach of fine-tuning the whole network. All experiments are performed on the CIFAR-100 dataset distributed across $10$ clients.
We argue that prompting is a more advantageous approach for adapting pre-trained models in Federated Learning, as it constrains the bias to the last layer instead of distributing it across the whole network. To experimentally prove this claim, we position ourselves at the end of the local training prior to the first synchronization with the server: namely, when each client is trained exclusively on its local distribution. In this scenario, we assess the Federated Bias within the feature space – which we refer to as feature bias for short – by computing the local prototypes for the $10$ clients and measuring their average pairwise Euclidean distance. Figure 3 (left) shows that prompting produces smaller feature bias compared to traditional fine-tuning, suggesting that the obtained features are less biased towards the local clients’ distributions. To validate this assertion, we also compute the same metric on the pre-trained model (ViT-B/16, on ImageNet-21k), which yields the smallest bias by design. This further supports our reasoning, as this last network experiences no fine-tuning whatsoever.
On the right-hand side of Figure 3, we show the performance of prompting vs. traditional fine-tuning. In line with recent works zhang2023slca ; panos2023first ; mcdonnell2024ranpac , prompting (Prompt) shows inferior performance compared to fine-tuning (FT), due to its limited plasticity. However, when leveraging Classifier Rebalancing (CR), we observe a reversal in this trend. Simply addressing Federated Bias in the last layer is sufficient for prompting to outperform classical fine-tuning: this suggests that prompting techniques train more biased classifiers while producing less biased features, effectively restricting FB to the final layer.
@@ 5 Relation with prior works
In its most naive form, Federated Learning combines local models at the conclusion of each round through element-wise parameters averaging mcmahan2017communication . Other approaches introduce regularization techniques during local models’ training to prevent them from drifting too far from the centralized parameter space of the server. For instance, FedProx li2020federated and SCAFFOLD karimireddy2020scaffold explicitly enforce such regularization. FedDC gao2022feddc estimates the local parameter shift and employs it as a correction term before aggregating parameters. GradMA luo2023gradma addresses optimization challenges (i.e., quadratic programming) by rerouting each client update towards optimizing the local problem while maintaining proximity to the server. On a different trajectory, the authors of luo2021no fit a Gaussian distribution for each class and use them to generate an IID dataset of features to calibrate the server’s classification head. FedProto tan2022fedproto also computes prototypes for each class observed by each client but aggregates them into global prototypes; these serve as target representations for subsequent rounds. In contrast, our approach samples from a hierarchical Gaussian Mixture Model, and incorporates prompting to effectively isolate Federated Bias and Incremental Bias to the last layer.
@@
Class-Incremental Learning (CIL) stands out as one of the most challenging settings within the Continual Learning domain vandeVen2022ThreeTO . In this scenario, the training process is divided into tasks, each introducing new classes as the training progresses. To tackle this setting, early techniques rely on regularization, establishing checkpoints for previous tasks to maintain proximity w.r.t. their parameters zenke2017continual ; kirkpatrick2017overcoming or predictions li2017learning . Alternatively, rehearsal-based approaches involve storing samples in a limited memory buffer and subsequently replaying them to optimize either the original objective robins1995catastrophic ; chaudhry2018efficient or a surrogate one based on knowledge distillation rebuffi2017icarl ; buzzega2020dark .
Lately, the emergence of pre-trained self-attentive architectures dosovitskiy2020image in the Computer Vision domain has paved the way for significant advancements, particularly with the advent of Parameter-Efficient Fine-Tuning (PEFT) techniques. Recent approaches wang2022learning ; wang2022dualprompt ; smith2023coda have harnessed prompting to achieve state-of-the-art performance in CIL, eliminating the need for a buffer to replay old samples. Instead, they employ a prompt pool comprising incrementally learned prompts, utilized to condition the network during the forward pass. In our approach, we adopt the prompting paradigm and introduce an efficient methodology that minimizes the computational overhead by eliminating the need for a double forward pass, ensuring state-of-the-art performance at the same time.
@@
The concept of Federated Class-Incremental Learning (FCIL) was initially introduced in yoon2021federated . In their work, the authors propose FedWeIT, which partitions client-side parameters into task-generic and task-specific components. To mitigate interference between clients, they implement sparse learnable masks to selectively extract relevant knowledge for each client. GLFC dong2022federated takes a different approach by combining local buffers with class-aware gradient compensation loss. This strategy helps counteract catastrophic forgetting through rehearsal, while also adjusting the magnitude of gradient updates based on whether input samples belong to new or old classes. Building upon this framework, an enhanced version is introduced by the same authors in the LGA paper dong2023no . TARGET zhang2023target tackles forgetting by training a centralized generator network to produce synthetic data, maintaining a similar behavior to the generator used in previous tasks. The generative network populates a buffer after each task, allowing clients to utilize it for rehearsal. Recent advancements in FCIL involve fine-tuning pre-trained models using PEFT techniques. Fed-CPrompt bagwe2023fed introduces a regularization term that encourages local prompts to diverge from global ones, enabling them to learn task-specific features. Instead, PILoRA guo2024federated integrates LoRA hu2021lora with prototypes, which are aggregated via a re-weighting mechanism on the server side during each communication round, whereas LoRM salami2024closed proposes a closed-form solution for merging parameter-efficient clients at the server side.